You want to take out a $243,000 mortgage (home loan). The interest rate on the loan is 5.6%, and the loan is for 30 years. Your monthly payments are $1,395.01. How much will still be owed after making payments for 5 years? Round your answer to the nearest dollar.
Answer #1

Interest rate per month = 5.6%/12

Total payments required = 30*12 = 360

Payments made = 5*12 = 60

Remaining balance in the loan account = A(P/A, i%, N – n)

= 1,395.01*(P/A, 5.6%, 360 - 60)

= 1,395.01*161.2711

= $224,974.86

The answer is $224,975